The thicknesses of 96 96 randomly selected linoleum tiles were found to have a variance of 3.07 3.07 . Construct the 95% 95 % confidence interval for the population variance of the thicknesses of all linoleum tiles in this factory. Round your answers to two decimal places.
Answer #1

n = 96

s2 = 3.07

= 0.05

ndf =n - 1 = 96 - 1 = 95

Confidence Interval:

Low End:

From Table:

Low End:

High End:

From Table:

High End:

So,

Confidence Interval:

2.35 < < 4.17